A point charge of 5.1 μC is located at the center of a sphere with a radius of 7 cm. The Coulomb constant is 8.98755 × 109 Nm2/C2, and the acceleration of gravity is 9.8 m/s2.

Determine the electric flux through the surface of the sphere.

Answer in units of N/C.
Answer #1
Given:
Charge q = 5.1 *10-6 C
radius of the sphere r = (7 cm)(1m/100cm)
                                    = 0.07 m
coulomb constant K = 8.98755*109 Nm2/C2
The magnitude of electric field everywhere on the surface of the sphere is
E = kq/r2 ...........(1)
now the electric field is perpendicular to the spherical surface at all points on the surface. The electric
flux through the surface is EA.
Here A is the area of the surface of the sphere = 4πr2
Electric flux φ = EA
                          = (kq/r2 )( 4πr2 )
                          = 4πkq
now plug the values in the above relation
φ = 4*(3.14)( 8.98755*109 Nm2/C2)( 5.1 *10-6 C)
    = 575.7*103 Nm2/C